# InfAlign: Inference-aware language model alignment

### v2

@@ Abstract
Language model alignment has become a critical step in training modern generative language models. The goal of alignment is to finetune a reference model such that the win rate of a sample from the aligned model over a sample from the reference model is high, subject to a KL divergence constraint. Today, we are increasingly using inference-time algorithms (e.g., Best-of-N, controlled decoding, tree search) to decode from language models rather than standard sampling. However, the alignment objective does not capture such inference-time decoding procedures. We show that the existing alignment framework is sub-optimal in view of such inference-time methods. We then modify the alignment objective and propose a framework for inference-aware alignment (InfAlign). We prove that for any inference-time decoding algorithm, the optimal solution that optimizes the inference-time win rate of the aligned policy against the reference policy is the solution to the typical RLHF problem with a transformation of the reward. This motivates us to provide the KL-regularized calibrate-and-transform RL (CTRL) algorithm to solve this problem, which involves a reward calibration step and a KL-regularized reward maximization step with a transformation of the calibrated reward. We particularize our study to two important inference-time strategies: best-of-$N$ sampling and best-of-$N$ jailbreaking, where $N$ responses are sampled from the model and the one with the highest or lowest reward is selected. We propose specific transformations for these inference-time strategies and demonstrate that our framework offers significant improvements over existing state-of-the-art methods for language model alignment. Empirically, we outperform state-of-the-art methods that are designed without taking inference-time decoding into consideration by 8-12% and 4-9% on inference-time win rates over the Anthropic helpfulness and harmlessness dialog benchmark datasets.
@@ 1 Introduction
Aligning generative language models (LMs) through KL-regularized reinforcement learning (KL-RL) is a widely adopted framework for finetuning a generative language model to improve a reward (e.g., safety or quality). Solving KL-RL generally entails training a reward model, and then using an RL solver (Christiano et al., 2017; Stiennon et al., 2020; Ouyang et al., 2022). Other ways of solving KL-RL include variants of direct preference optimization (Rafailov et al., 2023; Azar et al., 2023), reward model distillation (Fisch et al., 2024), and Best-of-$N$ distillation (Yang et al., 2024; Gui et al., 2024; Amini et al., 2024; Sessa et al., 2024). The success of the KL-RL framework is typically measured through the win rate of the aligned model over the reference model for a given task, which captures how often a sample drawn from the aligned model wins against a sample drawn from the base model using a judge for that task.
@@
However, rarely is the aligned model used as is at inference time; instead an inference-time procedure is typically used to accomplish a task. For example, it is customary to perform one or more of the following simple or complex procedures at decoding: best-of-$N$ sampling  (Nakano et al., 2022; Beirami et al., 2024), best-of-$N$ jailbreaking (Hughes et al., 2024b), chain-of-thought reasoning (Wei et al., 2022; OpenAI, 2024), and self consistency (Wang et al., 2022a) (see Section 7 for more). This creates a discrepancy between the inference-time decoding procedure and the training KL-RL objective, which is to maximize the expectation of a function of reward (e.g., win rate) of a sample from the aligned model against that of the base model.
@@
In this paper, we address the following question: Given a known inference-time procedure, can we align a model to optimize for the inference-time win rate against the reference model, where inference-time win rate entails obtaining a response from each model through the inference-time procedure and counting which sample wins? While directly optimizing the inference-time win rate seems intractable, we show that the optimal solution is captured via a family of optimization objectives, which we call the inference-aware alignment (InfAlign) framework. We further prove that for a $\delta$-bound language model, where the likelihood of all outcomes are upper-bounded by $\delta,$ as $\delta\to 0,$ the solution could be obtained by solving KL-RL with a specific transformation of the reward (Lemma 1). Therefore, the challenge of optimizing for inference-time win rate can be captured by designing a reward transformation that is suited to the specific inference-time procedure, and solving KL-RL using existing optimization algorithms like PPO (Schulman et al., 2017).
We particularize the study of InfAlign to two simple yet popular inference-time strategies: Best-of-$N$ (BoN) sampling (Nakano et al., 2022; Beirami et al., 2024) and Best-of-$N$ jailbreaking (Hughes et al., 2024a), which we call Worst-of-$N$ (WoN) since the defender may assume that the attacker is choosing the worst outcome of $N$ samples. Despite their simplicity, BoN is extremely effective, and even known to be an effective procedure for inference-time alignment (Beirami et al., 2024; Gui et al., 2024; Mudgal et al., 2024). WoN is a popular safety evaluation strategy often adopted to evaluate the ability to jailbreak foundation models (Yohsua et al., 2024; Chao et al., 2023; Souly et al., 2024; Hughes et al., 2024b). Hence, there is both theoretical and practical motivation for training to optimize inference-time BoN and WoN win rates. For these procedures, we solve InfAlign for different values of $N$ in the limit of infinitely expressive language models. We also show that exponential reward transformation is almost optimal for these strategies. We further show that the win rate optimal solution to KL-RL without reward transformation leads to suboptimal policies for inference-time win rate.
@@
Motivated by our theoretical findings, we propose a practical solver for InfAlign, called Calibrate-and-Transform Reinforcement Learning (CTRL), which adopts a three-step approach: (1) we first calibrate the scores of the reward model with respect to responses sampled per-prompt by the reference model; (2) we then adopt a suitable transformation of the calibrated scores for a given inference-time procedure; and (3) we solve the KL-RL problem (e.g., using PPO).
@@
We apply CTRL to the Anthropic helpfulness and harmlessness dataset (Bai et al., 2022) for optimizing BoN helpfulness and WoN harmlessness with various values of $N.$ We show that the (close-to-optimal) exponential reward transformations that were derived theoretically on idealized distributions transfer to real-world tasks – outperforming various KL-RL solvers at inference-time win rate. Finally, we also show CTRL (without any reward transformation) is a strong baseline for optimizing the standard win rate vs KL trade-off against various methods.
Our contributions are as follows:
  • We propose a framework for inference-aware language model alignment (InfAlign) that captures the optimal solution to optimizing win rate in view of any well-behaving inference-time procedure. Theoretically, we show that this framework satisfies a coupled-transformed reward/policy optimization objective which lends itself to iterative optimization for a large class of inference-time procedures (see Section 3).
  • Motivated by our theoretical findings, we propose a practical solver, calibrate-and-transform RL (CTRL) for InfAlign, which consists of an offline reward calibration and transformation step prior to applying the regular KL-RL framework. For two popular inference-time procedures, namely optimizing Best-of-$N$ (BoN) sampling and preventing Best-of-$N$ jailbreaking (WoN), we derive the conditions for optimal theoretical solutions to alignment in a toy setting. Somewhat surprisingly, for BoN and WoN, the optimal solution is independent of the base policy and the reward model, which leads to an iterative optimization approach that can be computed offline (see Section 4 and Section 5).
  • Empirically, we show on Anthropic dialog helpfulness and harmlessness datasets that CTRL with identity reward transformation achieves competitive performance compared to a variety of SOTA methods for optimizing standard win rate (Gui et al., 2024; Amini et al., 2024; Sessa et al., 2024; Azar et al., 2023; Rafailov et al., 2023), and produces better inference-time win rate vs KL tradeoffs by 8-12% for BoN and 4-9% for WoN inference-time procedures respectively (see Section 6).
@@ 2 Problem setup
[…]
@@
Let $r:{\bm{\mathcal{X}}}\times{\bm{\mathcal{Y}}}\rightarrow\mathbb{R}$ be a reward function that assigns a scalar value to any (prompt, response) pair, e.g., a model trained side-by-side on human preferences. The reward determines the goodness of response ${\bm{y}}$ in context ${\bm{x}}$. The goal of language model alignment is to construct an aligned distribution $\pi$ that improves the reward of the response while being close to ${\pi_{\rm ref}}$. Below we introduce KL-regularized reinforcement learning (RL), a popular training-time alignment framework in the literature.
Definition 1 (KL-regularized RL framework).
Let $\beta>0$ be a regularization parameter, the KL-regularized RL problem aims to maximize the pointwise expected reward with a KL regularizer below: $\forall{\bm{x}},$
$\pi^{*}_{r,\beta}(\cdot\mid{\bm{x}})=\arg\max_{\pi}\left\{\mathbb{E}_{{\bm{y}}\sim\pi(\cdot\mid{\bm{x}})}\{r({\bm{x}},{\bm{y}})\}-\beta D_{\text{KL}}(\pi(\cdot\mid{\bm{x}})\|{\pi_{\rm ref}}(\cdot\mid{\bm{x}}))\right\}.$ (1)
When evaluating an aligned policy, a common measure to use is the win rate over the base policy ${\pi_{\rm ref}}$. Let
$w_{r}({\bm{y}},{\bm{z}}\mid{\bm{x}}):=\mathbbm{1}\left\{r({\bm{x}},{\bm{y}})>r({\bm{x}},{\bm{z}})\right\}+\frac{1}{2}\mathbbm{1}\left\{r({\bm{x}},{\bm{y}})=r({\bm{x}},{\bm{z}})\right\}$
be the win random variable under reward $r$.
@@
Let $\mathcal{C}_{r,\pi}({\bm{x}},{\bm{y}})$ be the calibrated reward  [^2] The definition is similar to the cumulative density function of the reward $r({\bm{x}},{\bm{y}})$ under policy $\pi$ except for how ties are decided. under policy $\pi$ defined below
$\mathcal{C}_{r,\pi}({\bm{x}},{\bm{y}}):=\mathbb{E}_{{\bm{z}}\sim\pi(\cdot\mid{\bm{x}})}w_{r}({\bm{y}},{\bm{z}}\mid{\bm{x}}).$ (2)
The win rate is defined as following:
Definition 3 (Win rate).
For any policy $\pi_{1}$ and $\pi_{2}$, we define win rate of policy $\pi_{1}$ over policy $\pi_{2}$ given prompt ${\bm{x}}$, as measured by reward $r$ as follows:
$W_{r}(\pi_{1}\succ\pi_{2}\mid{\bm{x}}):=\mathbb{E}_{{\bm{y}}\sim\pi_{1}(\cdot\mid{\bm{x}})}\mathcal{C}_{r,\pi_{2}}({\bm{x}},{\bm{y}}).$ (3)
In this paper, for simplicity of the presentation and analysis, we assume that the language model is continuous, whose outcomes have infinitesimally small probabilities. In other words, we assume ${\bm{\mathcal{Y}}}$ is a continuous set, which can be mapped to $[0,1]$ through a CDF inverse transformation (Rosenblatt, 1952), with the ordering determined by the reward of the outcomes from the smallest reward to the highest reward. We also assume that $r$ assigns distinct rewards to different ${\bm{y}}$’s for a given ${\bm{x}}$. These assumptions have been made in the past implicitly by Hilton & Gao (2022) to estimate the KL divergence of Best-of-$n$ policy, and by Gui et al. (2024) to estimate its win rate and characterize the KL divergence and win rate of optimal policy for win rate vs KL divergence tradeoffs. While these assumptions lead to approximations when analyzing real-world distributions, the results derived under them are reasonably tight when the actual likelihood of the language model outcomes are small (Beirami et al., 2024). Note that we don’t make any such assumptions when providing our algorithmic developments, and the experimental results. Under these assumptions, we define the quantile mapping, which is the reverse of the calibrated reward mapping $\mathcal{C}^{-1}$, satisfying $\forall u\in[0,1]$,
$\mathcal{C}^{-1}_{r,\pi,{\bm{x}}}(u)={\bm{y}}_{u,{\bm{x}}}\quad\text{where}\quad\mathcal{C}_{r,\pi}({\bm{y}}_{u,{\bm{x}}}\mid{\bm{x}})=u.$
Inference-time processing.
As mentioned earlier, in many cases we do not simply just sample from the language model. When decoding is done through an inference-time procedure, the obtained sample follows a transformed distribution that depends on the aligned policy and the inference-time procedure. For example, when the inference-time procedure is BoN from the base policy $\pi$, it can be shown that the final obtained sample is distributed proportional to $\pi({\bm{y}}\mid{\bm{x}})\mathcal{C}_{r,\pi}({\bm{x}},{\bm{y}})^{N-1}$ (Beirami et al., 2024, Lemma 1). In this work, we model inference-time processing as a procedure $T$ that maps $\pi$ to a distribution $T(\pi)$,
$\pi\stackrel{{\scriptstyle T}}{{\longrightarrow}}T(\pi).$
In these cases, it is customary to compare the models by considering the following inference-time win rate of the aligned policy $\pi$.
@@
Under inference-time processing $T$, the inference-time win rate of policy $\pi_{1}$ over $\pi_{2}$ is defined as
$W^{T}_{r}(\pi_{1}\succ\pi_{2}\mid{\bm{x}}):=\mathbb{E}_{{\bm{y}}\sim T(\pi_{1})(\cdot\mid{\bm{x}}),{\bm{z}}\sim T(\pi_{2})(\cdot\mid{\bm{x}})}\left\{w_{r}({\bm{y}},{\bm{z}}\mid{\bm{x}})\right\}.$ (4)
With the above definitions at hand, the goal of the paper is solve the following KL-regularized inference-time win rate maximization problem.
Definition 5 (Inference-time win rate KL-regularized RL problem).
Let $T$ be a given inference-time procedure and $\beta>0.$ Then, the optimization problem for maximizing inference-time win rate is posed as
$\max_{\pi}\left\{W^{T}_{r}(\pi\succ{\pi_{\rm ref}}\mid{\bm{x}})-\beta D_{\text{KL}}(\pi(\cdot\mid{\bm{x}})\|{\pi_{\rm ref}}(\cdot|{\bm{x}}))\right\}.$ (5)
Note that this formulation reduces to the IPO objective (Azar et al., 2023, Equation (8)) when $T$ is the identity transformation and extends it otherwise for an arbitrary inference-time procedure. [^3] One question that arises is the role of the KL divergence regularizer in Eq. 5. We argue that the regularizer essentially enables multi-tasking between the SFT task and the RL task, which we formally prove for log-linear models in Appendix B. In other words, the KL divergence regularizer enables to preserve the core capabilities of the model while acquiring a new one through the KL-RL process.

### v5

@@ Abstract
Language model alignment is a critical step in training modern generative language models. Alignment targets to improve win rate of a sample from the aligned model against the base model. Today, we are increasingly using inference-time algorithms (e.g., Best-of-$N$, controlled decoding, tree search) to decode from language models rather than standard sampling. We show that this train/test mismatch makes standard RL framework sub-optimal in view of such inference-time methods. To this end, we propose a framework for inference-aware alignment (InfAlign), which aims to optimize inference-time win rate of the aligned policy against the base model. We prove that for any inference-time decoding procedure, the optimal aligned policy is the solution to the standard RLHF problem with a transformation of the reward. This motivates us to provide the calibrate-and-transform RL (InfAlign-CTRL) algorithm to solve this problem, which involves a reward calibration step and a KL-regularized reward maximization step with a transformation of the calibrated reward. For best-of-$N$ sampling and best-of-$N$ jailbreaking, we propose specific transformations offering up to 3-8% improvement on inference-time win rates. Finally, we also show that our proposed reward calibration method is a strong baseline for optimizing standard win rate.
@@ 1 Introduction
Aligning language models (LMs) through reinforcement learning from human feedback (RLHF) is a widely adopted finetuning framework to improve a reward (e.g., safety or quality). RLHF generally entails training a reward model, and then solving a KL-regularized RL problem (Christiano et al., 2017; Stiennon et al., 2020; Ouyang et al., 2022). Other ways of solving RLHF include variants of direct preference optimization (Rafailov et al., 2023; Azar et al., 2023), reward model distillation (Fisch et al., 2024), and Best-of-$N$ distillation (Gui et al., 2024; Amini et al., 2024; Sessa et al., 2024). The success of RLHF is typically measured through the win rate of the aligned model against the base model, i.e., how often a sample from the aligned model wins against the base model using a judge for the task.
@@
However, rarely is the aligned model used as is at inference time; instead an inference-time procedure is typically used to accomplish a task. For example, it is customary to perform one or more of the following procedures at decoding: best-of-$N$ sampling (Nakano et al., 2022; Beirami et al., 2024), best-of-$N$ jailbreaking (Hughes et al., 2024b), chain-of-thought reasoning (Wei et al., 2022; OpenAI, 2024), and self consistency (Wang et al., 2022a; DeepSeek-AI, 2025) (see Appendix A for a more comprehensive discussion).
[…]
$\mathcal{T}:$ Inference-time procedure
Figure 1: Given an inference-time procedure such as Best-of-NN, standard RLHF suffers from a train/test mismatch between train-time policy $\pi$ and inference-time policy $\mathcal{T}_{\pi}.$ InfAlign bridges the gap by optimizing a policy-transformed reward $\mathcal{R}$, yielding a policy, $\pi^{*},$ that is optimized for inference under $\mathcal{T}_{\pi^{*}}$.
@@
In this paper, we address the following question: Can we better align a language model to be served with a known inference-time procedure? We propose a family of optimization objectives, which we call the inference-aware alignment (InfAlign) framework. The objective optimizes for the inference-time win rate against the base model with KL regularization, where inference-time win rate entails obtaining a response from each model through the inference-time procedure and counting which sample wins. While directly optimizing the inference-time win rate seems intractable, we prove that the solution could be obtained by solving RLHF with a specific transformation of the reward (Lemma 1). Therefore, the challenge of optimizing for inference-time win rate can be captured by designing a reward transformation that is suited to the specific inference-time procedure. We present a diagram of InfAlign in Fig. 1.
We show that the optimal reward transformation satisfies a coupled-transformed reward/policy optimization objective which lends itself to iterative optimization for a large class of inference-time procedures (Theorem 1). However, the approach is unfortunately computationally inefficient and infeasible for real-world models.
@@
To enable practical solutions to InfAlign, we propose Calibrate-and-Transform Reinforcement Learning (InfAlign-CTRL), which adopts a three-step approach: (1) calibrate the scores of the reward model with respect to responses sampled per-prompt by the reference model; (2) transform the calibrated scores for a given inference-time procedure; and (3) solve the RLHF problem with the transformed reward. We prove various desirable properties of the reward calibration step and empirically show that it reduces reward hacking and improves standard win rate. Reward transformation allows us to further tailor the objective based on the inference-time procedure. Different choices of transformation lead to popular existing alignment objectives such as IPO (Azar et al., 2023) and best-of-$N$ distillation (Gui et al., 2024; Amini et al., 2024; Sessa et al., 2024).
We particularize the study to two simple yet popular inference-time strategies: Best-of-$N$ (BoN) sampling (Nakano et al., 2022; Beirami et al., 2024) and Best-of-$N$ jailbreaking (Hughes et al., 2024a), which we call Worst-of-$N$ (WoN) since the defender may assume that the attacker is choosing the worst outcome of $N$ samples. Despite simplicity, BoN is known to be an effective procedure for inference-time alignment (Beirami et al., 2024; Gui et al., 2024; Mudgal et al., 2024) and is the dominant approach in scaling inference-time compute (Snell et al., 2024; Brown et al., 2024). Variants of WoN are effective and popular for evaluating safety against jailbreaks (Yohsua et al., 2024; Chao et al., 2023; Souly et al., 2024; Hughes et al., 2024b; Beetham et al., 2024; Mehrabi et al., 2023). We find suitable reward transformations for these procedures through the InfAlign framework.
@@
Empirically, we apply InfAlign-CTRL to the Anthropic helpfulness (Bai et al., 2022), and Reddit summarization dataset (Stiennon et al., 2020) for optimizing BoN performance and Anthropic harmlessness for optimizing WoN performance with various values of $N.$ We show that solving InfAlign through InfAlign-CTRL outperforms various SOTA RLHF solvers at inference-time win rate by 3-8%. We also show InfAlign-CTRL (with no reward transformation) is on-par with SOTA methods for optimizing the standard win rate, with improved win rate for Anthropic helpfulness and harmlessness tasks.
Organization. In Section 2, we provide the InfAlign problem setup. In Section 3, we introduce RLHF with reward transformation as the main framework to solve InfAlign. In Section 4, we present the InfAlign-CTRL method, discuss properties of calibration, and present practical implementations. In Section 5, we provide experimental results.
@@ 2 Problem setup
We consider a generative language model that produces a response conditioned on an input prompt. Given a prompt $\bm{x}$}, e.g., ${\bm{x}}=\textit{What is a large language model?},$ a generative language model $\pi$, or a policy, specifies a conditional distribution $\pi(\cdot\mid{\bm{x}})$, from which the response $\bm{y}$} should be generated. We use $\bm{\mathcal{X}}$} and $\bm{\mathcal{Y}}$} to denote the space of possible inputs and outputs, respectively. Throughout the paper, we assume $\pi_{\rm ref}$} is a fixed base policy, e.g., obtained from supervised finetuning. We will often use the notation $\pi({\bm{y}}\mid{\bm{x}})\propto f({\bm{y}})$ for some $f:{\bm{\mathcal{Y}}}\rightarrow\mathbb{R}_{+}$ to denote the conditional distribution $\pi({\bm{y}}\mid{\bm{x}})=f({\bm{y}})/(\sum_{{\bm{y}}}f({\bm{y}}))$ obtained after normalization.
@@
Let $r:{\bm{\mathcal{X}}}\times{\bm{\mathcal{Y}}}\rightarrow\mathbb{R}$ be a reward function that assigns a scalar value to any (prompt, response) pair, e.g., a model trained side-by-side on human preferences. The reward determines the goodness of response $\bm{y}$} in context $\bm{x}$}. The goal of language model alignment is to construct an aligned distribution $\pi$ that improves the reward of the response while being close to $\pi_{\rm ref}$}. We introduce the popular KL-regularized reinforcement learning (RL) framework, also known as RLHF.
Definition 1 (KL-regularized RL).
Let $\beta>0$ be a regularization parameter, the KL-regularized RL (RLHF) problem aims to maximize the expected reward with a KL regularizer below: [^1] The solution to this optimization problem is unique and admits a closed-form expression (Korbak et al., 2022; Rafailov et al., 2023; Yang et al., 2024).
$\pi^{*}_{r,\beta}(\cdot\mid{\bm{x}})=\arg\max_{\pi}\left\{\mathcal{L}_{{\pi_{\rm ref}},r,\beta}\right\},\vskip-21.68121pt$
where
$\mathcal{L}_{{\pi_{\rm ref}},r,\beta}\!\!=\!\!\mathbb{E}_{{\bm{y}}\sim\pi(\cdot|{\bm{x}})}\{r({\bm{x}},{\bm{y}})\}\!\!-\!\!\beta D_{\text{KL}}(\pi(\cdot|{\bm{x}})\|{\pi_{\rm ref}}(\cdot|{\bm{x}})).$
When evaluating an aligned policy, a common measure to use is the win rate (Stiennon et al., 2020; Hilton & Gao, 2022) over the base policy $\pi_{\rm ref}$}. Let
$w_{r}({\bm{y}},{\bm{z}}|{\bm{x}})\!:=\!\!\mathbbm{1}\!\left\{r({\bm{x}},{\bm{y}})\!\!>\!\!r({\bm{x}},{\bm{z}})\right\}\!+\!\frac{\mathbbm{1}\!\left\{r({\bm{x}},{\bm{y}})\!\!=\!\!r({\bm{x}},{\bm{z}})\right\}}{2}$
be the win random variable under reward $r$. We define the calibrated reward and win rate below.
@@
The calibrated reward  [^2] The definition is similar to the cumulative density function (CDF) of the reward $r({\bm{x}},{\bm{y}})$ under $\pi$ except for how ties are decided. $\mathcal{C}_{r,\pi}({\bm{x}},{\bm{y}})$ under policy $\pi$ is defined below
$\mathcal{C}_{r,\pi}({\bm{x}},{\bm{y}}):=\mathbb{E}_{{\bm{z}}\sim\pi(\cdot\mid{\bm{x}})}\{w_{r}({\bm{y}},{\bm{z}}\mid{\bm{x}})\}.\vskip-14.45377pt$ (1)
Definition 3 (Standard win rate).
For any policy $\pi_{1}$ and $\pi_{2}$, the standard win rate (or win rate) of policy $\pi_{1}$ over policy $\pi_{2}$ given prompt $\bm{x}$}, as measured by reward $r$ is defined as:
$W_{r}(\pi_{1}\succ\pi_{2}\mid{\bm{x}}):=\mathbb{E}_{{\bm{y}}\sim\pi_{1}(\cdot\mid{\bm{x}}),{\bm{z}}\sim\pi_{2}(\cdot\mid{\bm{x}})}\{w_{r}({\bm{y}},{\bm{z}}\mid{\bm{x}})\}.$
Moreover, it can be shown that $W_{r}(\pi_{1}\;\succ\;\pi_{2}|{\bm{x}})=\mathbb{E}_{{\bm{y}}\sim\pi_{1}(\cdot|{\bm{x}})}\{\mathcal{C}_{r,\pi_{2}}({\bm{x}},{\bm{y}})\}.$
Inference-time procedure ($\mathcal{T}$).
As mentioned earlier, in many cases, decoding is done through an inference-time procedure. The obtained sample follows a transformed distribution that depends on the policy and the inference-time procedure. Let $\Delta_{\bm{\mathcal{Y}}}$ be the set of possible distributions over the output set $\bm{\mathcal{Y}}$}. In this work, we model an inference-time procedure as a mapping between distributions over the output space $\mathcal{T}:\Delta_{{\bm{\mathcal{Y}}}}\rightarrow\Delta_{{\bm{\mathcal{Y}}}}$,
$\pi(\cdot\mid{\bm{x}})\stackrel{{\scriptstyle\mathcal{T}}}{{\longrightarrow}}\mathcal{T}_{\pi}(\cdot\mid{\bm{x}}),\vskip-7.22743pt$
where $\mathcal{T}_{\pi}(\cdot\mid{\bm{x}})$ denotes the distribution of responses conditioned on $\bm{x}$} after the inference-time procedures is applied. In these cases, it is customary to compare the models by considering the following inference-time win rate of the aligned policy $\pi$.
@@
Under inference-time processing $\mathcal{T}$, the inference-time win rate of policy $\pi_{1}$ over $\pi_{2}$ is defined as
$W^{\mathcal{T}}_{r}(\pi_{1}\!\succ\!\pi_{2}|{\bm{x}})\!:=\!\mathbb{E}_{{\bm{y}}\sim\mathcal{T}_{\pi_{1}}(\cdot|{\bm{x}}),{\bm{z}}\sim\mathcal{T}_{\pi_{2}}(\cdot|{\bm{x}})}\{w_{r}({\bm{y}},{\bm{z}}|{\bm{x}})\}.\vskip-7.22743pt$
Similarly, it can be shown that
$W^{\mathcal{T}}_{r}(\pi_{1}\succ\pi_{2}\mid{\bm{x}})=\sum_{{\bm{y}}}\mathcal{C}_{r,\mathcal{T}_{\pi_{2}}}({\bm{x}},{\bm{y}})\mathcal{T}_{\pi_{1}}({\bm{y}}\mid{\bm{x}}),\vskip-10.84006pt$ (2)
where $\mathcal{C}_{r,\mathcal{T}_{\pi_{2}}}({\bm{x}},{\bm{y}})$ is the calibrated reward under the inference-time policy $\mathcal{T}_{\pi_{2}}$ (see Proof in Section B.2).
With the above definitions at hand, the goal of InfAlign is solve the following KL-regularized inference-time win rate maximization problem.
Definition 5 (InfAlign).
For a given inference-time procedure $\mathcal{T}$ and $\beta\!>\!0,$ InfAlign solves the following KL-regularized inference-time win rate maxmization problem:
$\max_{\pi}\hskip-1.4457pt\left\{W^{\mathcal{T}}_{r}(\pi\hskip-2.168pt\succ\hskip-2.168pt{\pi_{\rm ref}}|{\bm{x}})-\beta D_{\text{KL}}(\pi(\cdot|{\bm{x}})\|{\pi_{\rm ref}}(\cdot|{\bm{x}}))\right\}\hskip-2.168pt.$ (3)
The formulation of optimizing standard win-rate vs KL tradeoff has been previous studied for RLHF in Gui et al. (2024); Azar et al. (2023). The above formulation reduces to the IPO objective (Azar et al., 2023, Equation (8)) when $\mathcal{T}$ is the identity transformation and extends it otherwise for an arbitrary inference-time procedure. [^3] One question that arises is the role of the KL divergence regularizer in Eq. 3. We argue that the regularizer essentially enables multi-tasking between the SFT task and the RL task, which we formally prove for log-linear models in Appendix C. In other words, the KL divergence regularizer enables to preserve/distill the core capabilities of the SFT model while acquiring a new one through the RLHF process. In practice, the win rate is often evaluated by a judge different from the training time reward $r$. In this paper, we stick to $r$ when developing the algorithms as in standard RLHF framework (Definition 1). In experiments, we use a separate judge from the reward model.
Continuous language models.
For simplicity of the presentation and analysis, some of the theoretical results will be based on the assumption that $\bm{\mathcal{Y}}$} is a continuous set, and the language model has a density over $\bm{\mathcal{Y}}$}. We also assume that $r$ assigns distinct rewards to different $\bm{y}$}’s for a given $\bm{x}$}. Note that we don’t make any such assumptions when providing our algorithmic developments, and the experimental results.
These assumptions have been made in the past implicitly by (Hilton & Gao, 2022) to estimate the KL divergence of Best-of-$n$ policy, and by (Gui et al., 2024) to characterize the KL divergence and win rate tradeoffs. While these assumptions lead to approximations when analyzing real-world distributions, the results derived under them are reasonably tight when the actual likelihood of the language model outcomes are small (Beirami et al., 2024).